OpenStudy (epicteatime):

(4,1); perpendicular to y=1/3x+3. How to solve it?

OpenStudy (epicteatime):

@Jadedry

OpenStudy (epicteatime):

@Hero @ganeshie8

OpenStudy (epicteatime):

@ParthKohli @mathstudent55

ganeshie8 (ganeshie8):

What exactly are you trying to solve here ?

OpenStudy (epicteatime):

I want to know the method to solve the problem

OpenStudy (epicteatime):

I need to use like y-y1=m(x-x1)

OpenStudy (epicteatime):

to find the equation

ganeshie8 (ganeshie8):

that's fine, may i know what equation you're trying to find ?

ganeshie8 (ganeshie8):

the problem is not clear to me yet..

OpenStudy (jadedry):

@Epicteatime Hi again! I too am a bit confused as to what you are trying to find. :c

OpenStudy (epicteatime):

the equation for the line that passes through the given point and satisfies the given condition. And Hi jade :3

OpenStudy (epicteatime):

Like, do I use the opposite slope? -3/1 since its perpendicular?

OpenStudy (jadedry):

So the line has to be perpendicular to y = 1 and it has to pass through the point 4,1 ?

OpenStudy (epicteatime):

yes, basically

OpenStudy (anonymous):

you have an X and the Y just plug in the numbers and then get the answer of Y = 1/15

OpenStudy (jadedry):

Then the equation for the line should be x = 4 It would be perpendicular to you, which is a horizontal line, and it would pass through 4,1

OpenStudy (jadedry):

Unless there is another piece of info I should take into account?

OpenStudy (epicteatime):

hmmm

OpenStudy (epicteatime):

like if it was parallel, I could have used the slope from the line stuff that is given.

OpenStudy (jadedry):

Wait, Is the equation for Y \[y = \frac{ 1 }{ 3x+3 }\] ?

OpenStudy (epicteatime):

uhh, I have no idea xD Thats why I'm here xD

OpenStudy (jadedry):

Your textbook should tell you though, the format in which you wrote it was a bit confusing, I thought y=1 at first?

OpenStudy (epicteatime):

hmm, maybe maybe xD

OpenStudy (epicteatime):

its ok, i'll ask my teacher/ look at my textbooks :3 gnat :D

OpenStudy (epicteatime):

*gnyt

OpenStudy (jadedry):

Ah, okay. Good night! c:

OpenStudy (anonymous):

Y does equal 1 it says so in the point for the graph( 4,1)

OpenStudy (epicteatime):

yep, completely :D